# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 1.2. Rigidity and existence of discrete conformal structures
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
@@
For the discrete conformal structure (A1), let $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ be the weights on $(\Sigma,\mathcal{T})$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ if $\alpha_{i}=\alpha_{j}$ for any two adjacent boundary components $i,j\in B$;
@@
For the discrete conformal structure (A2), let $\eta\in\mathbb{R}_{>0}^{E}$ be the weight on $(\Sigma,\mathcal{T})$;
@@
For the mixed discrete conformal structure $\mathrm{I}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}$ are induced by (7) and (10), let $\xi^{(1)}\in\mathbb{R}_{>0}^{E}$ be the weights on $(\Sigma,\mathcal{T})$;
@@
For the mixed discrete conformal structure $\mathrm{II}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}$ are induced by (5) and (9), let $\alpha:B\rightarrow\{-1,0,1\}$ and $\xi^{(2)}:E\rightarrow(1,+\infty)$ be the weights on $(\Sigma,\mathcal{T})$;
Then the discrete conformal factor $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. In particular, the map from $f$ to $K$ is a smooth embedding. Furthermore, the image of $K$ is >$\mathbb{R}^{N}_{>0}$ for the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$, for the case $\mathrm{(i)}$ if $\alpha:B\rightarrow\{0,1\}$, and for the case $\mathrm{(iv)}$ if $\alpha:B\rightarrow\{0,1\}$ and satisfies a restrictive condition.
@@
If $\alpha\equiv 0$, the case $\mathrm{(i)}$ in Theorem 1.9 generalizes Guo’s results [4] and Li-Xu-Zhou’s results [6]. If $\alpha\equiv 1$, the case $\mathrm{(i)}$ in Theorem 1.9 was proved by Guo-Luo [5]. Moreover, the case $\mathrm{(ii)}$ in Theorem 1.9 was also proved by Guo-Luo [5]. We include them in Theorem 1.9 for completeness. For the case $\mathrm{(iii)}$ in Theorem 1.9, the weight $\xi^{(1)}\in\mathbb{R}_{>0}^{E}$ can be extended to a wider range, please see Theorem 5.1. For the case $\mathrm{(iv)}$ in Theorem 1.9, the weight $\xi^{(2)}:E\rightarrow(1,+\infty)$ can also be extended to a wider range, and we will use a new notation $\Xi$ to denote the set of the range of weight because of its complexity, please see Subsection 6.1. Moreover, for the image of $K$ in the mixed discrete conformal structure $\mathrm{II}$, we not only require $\alpha:B\rightarrow\{0,1\}$, but also $\alpha$ satisfies a restrictive condition. Here the restrictive condition means to exclude the case that $\alpha_{j}=\alpha_{k}=1$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with the edge lengths $l_{ij}$ and $l_{ik}$ are given by (9) and the edge length $l_{jk}$ is given by (5), please see Subsection 6.3.
@@
The proof for the existence of discrete conformal structures uses the continuity method. We will show that the image of the curvature map is both open and closed subset of >$\mathbb{R}^{N}_{>0}$, and thus is the whole space >$\mathbb{R}^{N}_{>0}$. The proof for the rigidity of discrete conformal structures involves a variational principle, which was pioneered by Colin de Verdière’s work [1] for tangential circle packings on triangulated surfaces. The proof of the rigidity consists of the following three steps.
The first step is to give a characterization of the admissible space of discrete conformal factor $f$ for a right-angled hyperbolic hexagon. By a change of variables $u_{i}=u_{i}(f_{i})$ described in the proof of Lemma 2.22, the admissible space $\Omega_{ijk}$ of $f$ is transferred to the admissible space $\mathcal{U}_{ijk}$ of $u$, which is proved to be convex and simply connected. The second step is to prove the Jacobian of the generalized angles $\theta$ with respect to $u$ for a right-angled hyperbolic hexagon is symmetric and negative definite. As a consequence, the Jacobian of the generalized combinatorial curvature $K$ with respect to $u$ is symmetric and negative definite on the admissible space $\mathcal{U}=\cap\ \mathcal{U}_{ijk}$. The first step and second step enable us to define a strictly concave function on the admissible space $\mathcal{U}$. The third step is to use the following well-known result from analysis:
@@
The key step is how to get a change of variables $u_{i}=u_{i}(f_{i})$. For this, we need to calculate the variation formulas of the generalized angles, which is as follows.
@@
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
@@
For any right-angled hyperbolic hexagon $\{ijk\}\in F$,
@@
where $\beta=1$ if $c_{ijk}$ is time-like and $\beta=-1$ if $c_{ijk}$ is space-like. If $c_{ijk}$ is light-like, we interpret the formula as $\tanh^{\beta}h_{ij}=\tanh^{\beta}h_{jk}=\tanh^{\beta}h_{ki}=1$.
@@
Furthermore, we have the following Glickenstein-Thomas formula for $\frac{\partial\theta^{jk}_{i}}{\partial f_{i}}$, $\frac{\partial\theta^{ik}_{j}}{\partial f_{i}}$ and $\frac{\partial\theta^{ij}_{k}}{\partial f_{i}}$, i.e.,
@@
Moreover, there is a change of variables $u_{i}=u_{i}(f_{i})$ such that
(11) $\frac{\partial\theta^{jk}_{i}}{\partial u_{j}}=\frac{\partial\theta^{ik}_{j}}{\partial u_{i}}.$
The function $u$ is also called a discrete conformal factor.
Here $h_{ij}$ is the signed distance of the face center $c_{ijk}$ to the hyperbolic geodesic $E_{ij}$, which is positive if $c_{ijk}$ is on the same side of the geodesic $E_{ij}$ as the right-angled hyperbolic hexagon $\{ijk\}$ and negative otherwise (or zero if $c_{ijk}$ is on the geodesic $E_{ij}$).
@@ 1.4. Organization of the paper
In Section 2, we prove Theorem 1.12 by some useful lemmas. In Section 3, we prove Theorem 1.9 (i). In Section 4, we prove Theorem 1.9 (ii). In Section 5, we prove Theorem 1.9 (iii). In Section 6, we prove Theorem 1.9 (iv).
@@ 2. Discrete conformal variations of generalized angles
In this section, we first present some basic facts in the 2-dimensional hyperbolic geometry, which are discussed in [7, Chapter 3 and 6]. Then we prove Theorem 1.12. Different from the proof of Theorem 7 in Glickenstein-Thomas [3], we are unable to get the variational formulas of generalized angle by similar geometric properties directly from the figure under the $f_{k}$-conformal variation, and we obtain them by complex calculations.
2.1. Some basic knowledge
For any vectors $x,y$ in $\mathbb{R}^{3}$, the Lorentzian inner product * of $x$ and $y$ is given by $x*y=xJy^{T}$, where $J=\mathrm{diag}\{1,1,-1\}$. If $x*y=0$, then we say $x,y\in\mathbb{R}^{3}$ are Lorentz orthogonal. The Lorentzian cross product $\otimes$ of $x$ and $y$ is defined as $x\otimes y:=J(x\times y)$, where $\times$ is the Euclidean cross product.
The Lorentzian norm of $x\in\mathbb{R}^{3}$ is defined to be the complex number $||x||=\sqrt{x*x}$. Here $||x||$ is either positive, zero or positive imaginary. If $||x||$ is positive imaginary, we denote by $|||x|||$ its absolute value. Furthermore, $x\in\mathbb{R}^{3}$ is called space-like if $x*x>0$, light-like if $x*x=0$ and time-like if $x*x<0$. The hyperboloid model of $\mathbb{H}^{2}$ is defined to be
$\mathbb{H}^{2}=\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}|\ x*x=-1\ \text{and}\ x_{3}>0\},$
which can be embedded in $(\mathbb{R}^{3},*)$. A geodesic (or hyperbolic line) in $\mathbb{H}^{2}$ is the non-empty intersection of $\mathbb{H}^{2}$ with a 2-dimensional vector subspace of $\mathbb{R}^{3}$.
The Klein model of $\mathbb{H}^{2}$ is the center projection of $\mathbb{H}^{2}$ into the plane $x_{3}=1$ of $\mathbb{R}^{3}$. In the Klein model, the interior, boundary and external points of the unit disk represent the time-like, light-like and space-like vectors respectively, and they are called hyperbolic points, ideal points and hyper-ideal points respectively. A geodesic in $\mathbb{H}^{2}$ is a chord on the unit disk. In the Klein model, a triangle $xyz$ with its three vertices $x,y,z$ being hyperbolic is called a hyperbolic triangle. If at least one of $x,y,z$ is ideal or hyper-ideal, then we say the triangle $xyz$ is a generalized hyperbolic triangle. Specially, if $x,y,z$ are all hyper-ideal, we call this triangle as a hyper-ideal hyperbolic triangle for simplicity.
The following proposition follows from the properties of Lorentzian cross product, which is used extensively in this paper.
@@ 2.1. Some basic knowledge
For two vectors $x,y\in\mathbb{H}^{2}$, we denoted by $d_{\mathbb{H}}(x,y)$ the hyperbolic distance between $x$ and $y$, which is the arc length of the geodesic between $x$ and $y$. We use $\mathrm{Span}(x)$ or $\mathrm{Span}(x,y)$ to denote the vector subspace spanned $x$ or $x,y$. If $x\in\mathbb{R}^{3}$ is a space-like vector, then the Lorentzian complement $x^{\bot}:=\{w\in\mathbb{R}^{3}|x*w=0\}$ of $\mathrm{Span}(x)$ is a 2-dimension vector subspace of $\mathbb{R}^{3}$ and $x^{\bot}\cap\mathbb{H}^{2}$ is non-empty. This implies that $x^{\bot}\cap\mathbb{H}^{2}$ is a geodesic in $\mathbb{H}^{2}$. And the hyperbolic distance between $x$ and $y\in\mathbb{H}^{2}$ is defined to be $d_{\mathbb{H}}(y,x^{\bot})=\inf\{d_{\mathbb{H}}(y,z)|z\in x^{\bot}\cap\mathbb{H}^{2}\}$. Let $x,y\in\mathbb{R}^{3}$ be space-like vectors. If $\mathrm{Span}(x,y)\cap\mathbb{H}^{2}$ is non-empty, then the hyperbolic distance between $x$ and $y$ is defined by $d_{\mathbb{H}}(x^{\bot},y^{\bot})=\inf\{d_{\mathbb{H}}(z,w)|z\in x^{\bot}\cap\mathbb{H}^{2},w\in y^{\bot}\cap\mathbb{H}^{2}\}$. If $\mathrm{Span}(x,y)\cap\mathbb{H}^{2}$ is empty, then $x^{\bot}$ and $y^{\bot}$ intersect in an angle $\angle(x^{\bot},y^{\bot})$ in $\mathbb{H}^{2}$.
@@
Suppose $x,y\in\mathbb{H}^{2}$ are time-like vectors and $z,w\in\mathbb{R}^{3}$ are space-like vectors. Then
@@
$x*y=-|||x|||\cdot|||y|||\cosh d_{\mathbb{H}}(x,y)$;
@@
$|x*z|=|||x|||\cdot||z||\sinh d_{\mathbb{H}}(x,z^{\bot})$, and $x*z<0$ if and only if $x$ and $z$ are on opposite sides of the hyperplane $z^{\bot}$;
@@
if $\mathrm{Span}(z,w)\cap\mathbb{H}^{2}$ is non-empty, then $|z*w|=||z||\cdot||w||\cosh d_{\mathbb{H}}(z^{\bot},w^{\bot})$, and $z*w<0$ if and only if $z$ and $w$ are oppositely oriented tangent vectors of $Q$, where $Q$ is the hyperbolic line Lorentz orthogonal to $z^{\bot}$ and $w^{\bot}$;
@@
if $\mathrm{Span}(z,w)\cap\mathbb{H}^{2}$ is empty, then $z*w=||z||\cdot||w||\cos\angle(z^{\bot},w^{\bot})$.
The following proposition is useful for identifying perpendicular geodesics in the Klein model.
@@
([8], Proposition 4) Let $x\in\mathbb{R}^{3}$ be a space-like vector and let $\gamma=x^{\bot}\cap\mathbb{H}^{2}$ be the geodesic corresponding to $x^{\bot}$. Suppose $\omega$ is another geodesic that intersects $\gamma$, and $P$ is the plane such that $\omega=P\cap\mathbb{H}^{2}$. Then $\gamma$ and $\omega$ meet at a right angle if and only if $x\in P$.

### v3

@@ 1.2. Main results
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary, and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
@@
For the discrete conformal structure (A1), let $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ be the weights on $(\Sigma,\mathcal{T})$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ for any two adjacent boundary components $i,j\in B$. Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. Furthermore, if $\alpha:B\rightarrow\{0,1\}$, then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
@@
For the discrete conformal structure (A2), let $\eta\in(-1,+\infty)^{E}$ be the weight on $(\Sigma,\mathcal{T})$. Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. Furthermore, if $\eta\in(-1,0]^{E}$, then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
@@
For the discrete conformal structure (A3), let $\eta\in\mathbb{R}_{>0}^{E}$ be the weight on $(\Sigma,\mathcal{T})$. Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. Furthermore, the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
@@
For the mixed discrete conformal structure I, let $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in(1,+\infty)^{E}$ be the weights on $(\Sigma,\mathcal{T})$. Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. Furthermore, if $\alpha:B\rightarrow\{0,1\}$, and $\alpha_{j}$ and $\alpha_{k}$ cannot both be 1 simultaneously for any right-angled hyperbolic hexagon with edge lengths $l_{ij},l_{ik}$ given by (8) and $l_{jk}$ given by (5), then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
(v):
For the mixed discrete conformal structure II, let $\eta\in[1,+\infty)^{E}$ be the weight on $(\Sigma,\mathcal{T})$. Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
(vi):
For the mixed discrete conformal structure III, let $\eta\in\mathbb{R}_{>0}^{E}$ be the weight on $(\Sigma,\mathcal{T})$. Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. Furthermore, the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
@@
If $\alpha\equiv 0$, then Theorem 1.8 (i) generalizes the results obtained by Guo [4] and Li-Xu-Zhou [6]. For $\alpha\equiv 1$, Theorem 1.8 (i) was established by Guo-Luo [5]. Theorem 1.8 (i) unifies these existing results and further incorporates mixed-type cases where $\alpha_{i}=1$ for some non-empty subset $V_{1}\subseteq V$ and $\alpha_{j}=0$ for $j\in V\setminus V_{1}$. Theorem 1.8 (ii) generalizes the results of Guo-Luo [5], while Theorem 1.8 (iii), which was proven by Guo-Luo [5], is included here for completeness. In Theorem 1.8 (iv), the weight $\eta\in(1,+\infty)^{E}$ admits an extension to a broader range, with details provided in Subsection 8.1. For Theorem 1.8 (v), the weight $\eta\in[1,+\infty)^{E}$ can also be extended to a broader range (see Theorem 7.5). Note that only a rigidity result is presented here, without any existence result; the reason is elaborated in Remark 7.6. In Theorem 1.8 (vi), the weight $\eta\in\mathbb{R}_{>0}^{E}$ similarly allows for extension to a broader range, as discussed in Theorem 6.1.
@@
The proof of the existence of discrete conformal structures relies on the continuity method. Specifically, we will show that the image of the curvature map $K$ is both an open and a closed subset of >$\mathbb{R}^{N}_{>0}$; consequently, it coincides with the entire space >$\mathbb{R}^{N}_{>0}$.
The proof of the rigidity of discrete conformal structures involves a variational principle introduced by Colin de Verdière [1] for tangential circle packings on triangulated closed surfaces. The rigidity proof proceeds in three steps: First, we characterize the admissible space of discrete conformal factors $f$ for a right-angled hyperbolic hexagon. By a variable transformation $u_{i}=u_{i}(f_{i})$ (as detailed in the proof of Lemma 2.21), the admissible space $\Omega_{ijk}$ of $f$ is transformed into the admissible space $\mathcal{U}_{ijk}$ of $u$, which is shown to be convex and simply connected. Second, we prove that the Jacobian of the generalized angles $\theta$ with respect to $u$ for a right-angled hyperbolic hexagon is symmetric and negative definite. Consequently, the Jacobian of the generalized combinatorial curvature $K$ with respect to $u$ is symmetric and negative definite on the admissible space $\mathcal{U}=\cap\,\mathcal{U}_{ijk}$. These two steps collectively enable the definition of a strictly concave function on $\mathcal{U}$. Third, we invoke the following well-known result from analysis to establish rigidity.
@@
A critical component of the proof involves constructing the necessary variable transformation $u_{i}=u_{i}(f_{i})$. To this end, we require the following result concerning the variation of generalized angles.
@@
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary, and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
@@
For any right-angled hyperbolic hexagon $\{ijk\}\in F$, the following holds:
@@
where $\beta=1$ if the face center $c_{ijk}$ is time-like, and $\beta=-1$ if $c_{ijk}$ is space-like. If $c_{ijk}$ is light-like, the formula is interpreted as $\tanh^{\beta}h_{ij}=\tanh^{\beta}h_{jk}=\tanh^{\beta}h_{ki}=1$. Here, $h_{ij}$ denotes the signed distance from the face center $c_{ijk}$ to the hyperbolic geodesic line $E_{ij}$, which is positive if $c_{ijk}$ lies on the same side of $E_{ij}$ as the right-angled hyperbolic hexagon $\{ijk\}$, negative otherwise, and zero if $c_{ijk}$ lies on $E_{ij}$.
@@
For a right-angled hyperbolic hexagon $\{ijk\}\in F$, the variation of generalized angles satisfies the following relation:
@@
There exists a change of variables given by $u_{i}=u_{i}(f_{i})$ such that
The function $u$ is also referred to as a discrete conformal factor.
@@ 1.4. Organization of the paper
In Section 2, we establish Theorem 1.11. In Section 3, we investigate the discrete conformal structure (A1) and prove Theorem 1.8 (i). Section 4 is devoted to the discrete conformal structure (A2), where Theorem 1.8 (ii) is proven. In Section 5, we study the discrete conformal structure (A3) and establish Theorem 1.8 (iii). Section 6 focuses on the mixed discrete conformal structure III, with a proof of a generalization of Theorem 1.8 (vi). Similarly, Section 7 investigates the mixed discrete conformal structure II and proves a generalization of Theorem 1.8 (v). Finally, Section 8 studies the mixed discrete conformal structure I and establishes a generalization of Theorem 1.8 (iv).
Acknowledgements
The research of X. Xu is supported by National Natural Science Foundation of China under grant no. 12471057.
@@ 2. Discrete conformal variations of generalized angles
In this section, we first recall some fundamental results from 2-dimensional hyperbolic geometry (see [7, Chapters 3 and 6]). We then provide a geometric interpretation along with several useful lemmas. Finally, we proceed to prove Theorem 1.11.
2.1. Some fundamental results
For any vectors $x,y$ in the Lorentzian space $\mathbb{R}^{3}$, the Lorentzian inner product * of $x$ and $y$ is defined as $x*y=xJy^{\mathrm{T}}$, where $J=\mathrm{diag}\{1,1,-1\}$. Two vectors $x,y\in\mathbb{R}^{3}$ are said to be Lorentz orthogonal if $x*y=0$. The Lorentzian cross product $\otimes$ of $x$ and $y$ is defined as $x\otimes y=J(x\times y)$, where $\times$ denotes the Euclidean cross product.
The Lorentzian norm of a vector $x\in\mathbb{R}^{3}$ is defined as the complex number $\|x\|=\sqrt{x*x}$. Here $\|x\|$ is either positive, zero, or positive imaginary. When $\|x\|$ is positive imaginary, its absolute value is denoted by $|\|x\||$. Furthermore, a vector $x\in\mathbb{R}^{3}$ is termed space-like if $x*x>0$, light-like if $x*x=0$, and time-like if $x*x<0$. The hyperboloid model of $\mathbb{H}^{2}$ is defined as
$\mathbb{H}^{2}=\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}\mid x*x=-1\ \text{and}\ x_{3}>0\},$
which is embedded in the Lorentzian space $(\mathbb{R}^{3},*)$. A vector subspace of $\mathbb{R}^{3}$ is termed time-like if it contains at least one time-like vector. A geodesic (or hyperbolic line) in $\mathbb{H}^{2}$ is the non-empty intersection of $\mathbb{H}^{2}$ with a 2-dimensional time-like vector subspace of $\mathbb{R}^{3}$.
The Klein model of $\mathbb{H}^{2}$ is defined as the central projection of $\mathbb{H}^{2}$ onto the plane $x_{3}=1$ in $\mathbb{R}^{3}$. In this model, the interior, boundary, and exterior points of the unit disk correspond to time-like, light-like, and space-like vectors, respectively, and are referred to as hyperbolic points, ideal points, and hyper-ideal points, correspondingly. In the Klein model, a geodesic in $\mathbb{H}^{2}$ is represented as an open chord of the unit disk. A triangle $xyz$ with all three vertices $x,y,z$ being hyperbolic is called a hyperbolic triangle. If at least one of $x,y,z$ is ideal or hyper-ideal, then the triangle $xyz$ is termed a generalized hyperbolic triangle. Specifically, if $x,y,z$ are hyper-ideal, this triangle is referred to as a hyper-ideal hyperbolic triangle for brevity. It should be noted that, in this paper, each edge of a generalized hyperbolic triangle $xyz$ is required to intersect the unit disk.
The following proposition, which follows from properties of the Lorentzian cross product, is used extensively in this paper.
@@ 2.1. Some fundamental results
For any two vectors $x,y\in\mathbb{H}^{2}$, there exists a unique geodesic passing through $x$ and $y$. The arc length of the geodesic segment between $x$ and $y$ is defined as the hyperbolic distance between $x$ and $y$, denoted by $d_{\mathbb{H}}(x,y)$. We use $\textrm{Span}(x)$ and $\textrm{Span}(x,y)$ to denote the vector subspace spanned by $x$ and by $x,y$, respectively. For a space-like vector $x\in\mathbb{R}^{3}$, the Lorentzian complement of $\textrm{Span}(x)$, defined as $x^{\bot}:=\{w\in\mathbb{R}^{3}\mid x*w=0\}$, is a 2-dimensional time-like vector subspace of $\mathbb{R}^{3}$, and thus $x^{\bot}\cap\mathbb{H}^{2}$ is a geodesic in $\mathbb{H}^{2}$. The hyperbolic distance between $x$ and $y\in\mathbb{H}^{2}$ is defined as $d_{\mathbb{H}}(y,x^{\bot})=\inf\{d_{\mathbb{H}}(y,z)\mid z\in x^{\bot}\cap\mathbb{H}^{2}\}$. Let $x,y\in\mathbb{R}^{3}$ be space-like vectors. If $\textrm{Span}(x,y)\cap\mathbb{H}^{2}$ is non-empty, then the hyperbolic distance between $x$ and $y$ is defined by $d_{\mathbb{H}}(x^{\bot},y^{\bot})=\inf\{d_{\mathbb{H}}(z,w)\mid z\in x^{\bot}\cap\mathbb{H}^{2},w\in y^{\bot}\cap\mathbb{H}^{2}\}$. If $\textrm{Span}(x,y)\cap\mathbb{H}^{2}$ is empty, then $x^{\bot}$ and $y^{\bot}$ intersect at an angle $\angle(x^{\bot},y^{\bot})$ in $\mathbb{H}^{2}$.
@@
Suppose $x,y\in\mathbb{H}^{2}$ are time-like vectors, and $z,w\in\mathbb{R}^{3}$ are space-like vectors. Then
@@
$x*y=-|\|x\||\cdot|\|y\||\cosh d_{\mathbb{H}}(x,y)$;
@@
$|x*z|=|\|x\||\cdot\|z\|\sinh d_{\mathbb{H}}(x,z^{\bot})$, and $x*z<0$ if and only if $x$ and $z$ lie on opposite sides of the hyperplane $z^{\bot}$;
@@
if $\textrm{Span}(z,w)\cap\mathbb{H}^{2}\neq\emptyset$, then $|z*w|=\|z\|\cdot\|w\|\cosh d_{\mathbb{H}}(z^{\bot},w^{\bot})$, and $z*w<0$ if and only if $z$ and $w$ are oppositely oriented tangent vectors of $Q$, where $Q$ is the hyperbolic line Lorentz orthogonal to $z^{\bot}$ and $w^{\bot}$;
@@
if $\textrm{Span}(z,w)\cap\mathbb{H}^{2}=\emptyset$, then $z*w=\|z\|\cdot\|w\|\cos\angle(z^{\bot},w^{\bot})$.
The following two propositions are useful for characterizing perpendicular geodesics in the Klein model.
@@
([7], Theorem 3.2.11) Let $P$ be a geodesic in $\mathbb{H}^{2}$, and let $y\in\mathbb{H}^{2}$ be a time-like vector. There exists a unique geodesic in $\mathbb{H}^{2}$ passing through $y$ and Lorentz orthogonal to $P$.